OpenStudy (anonymous):

how many moles of MgS2O3 are in 205 g of the compound?

OpenStudy (ookawaiioo):

MW: 136 grams of MgS2O3 per mole 206 g of MgS2O3 * 1mole/ 136 grams = 1.514 moles

OpenStudy (anonymous):

no of moles=mass / Mr = 206/136 =1.514

OpenStudy (ookawaiioo):

Jocking my answer.....

OpenStudy (anonymous):

@oOkawaiiOo..i have just answer ze question...wats zer to joke in ur answer..